# MoD-SLAM: Monocular Dense Mapping for Unbounded 3D Scene Reconstruction

### 26 Feb 2024 v4

@@ Abstract
Neural implicit representations have recently been demonstrated in many fields including Simultaneous Localization And Mapping (SLAM). Current NeRF-based SLAM excels in reconstructing bounded scenes with accurate depth and RGB image input. However, relying solely on RGB image input fails to reconstruct the scene’s scale accurately, and also leads to scale drift due to errors accumulated during tracking. To address these issues, we present MoD-SLAM, a monocular dense mapping method that allows global pose optimization and 3D reconstruction in real-time in unbounded scenes. Optimizing scene reconstruction by monocular depth estimation and using loop closure detection to update camera pose enable detailed and precise reconstruction on large scenes. Compared to recent works, our approach is more robust, scalable and versatile. Our experiments demonstrate that MoD-SLAM has state-of-the-art monocular mapping performance, especially in large borderless scenes.
@@
3D Scene Reconstruction, Neural Radiance Field, SLAM, Monocular Depth Estimation, Unbounded Scene
@@ 1 Introduction
In recent years, due to the development and advancement of Neural Radiance Field (NeRF) [15, 8] and view synthesis [19, 26, 1] in 3D scene reconstruction, dense visual SLAM mapping has become feasible. Early NeRF-based SLAM systems such as iMAP [30] and NICE-SLAM [46] have achieved ideal results in dense mapping and tracking. However, they all have the problems of decreasing accuracy and scale drift in large scenes. Subsequent researches [42, 43, 38, 6, 25] apply the signed distance function (SDF) [22] to reconstruct the 3D scene more accurately. Nevertheless, these approaches still present limited capability in unbounded scenes. Furthermore, numerous works rely on RGB-D input, which fails to adequately reconstruct the scale of the scene, the texture structure, and the details of geometric objects in the monocular mode. To address this challenge, GO-SLAM [43] has been conducted to adapt NeRF-based SLAM to function solely with RGB input. However, It lacks the prior depth maps and thus, cannot properly reconstruct the scale of the scene (Fig.1). In addition, numerous previous works lack a loop closure detection, which results in accumulated positional drift, leading to inaccuracies in tracking and mapping process.
@@
Figure 1: Reconstruction results on ScanNet 0207 [7]. We present MoD-SLAM, a neural-based monocular dense mapping method. Our method shows a more powerful capability than existing work: GO-SLAM [43].
In this work, we propose MoD-SLAM, a neural implicit dense visual SLAM system based on monocular input. Unlike previous NeRF-based SLAM systems, our method eliminates the strict requirement for RGB-D data. We improve its accuracy by incorporating a monocular depth estimation module and a depth distillation module at the backend. These two modules allow the system to generate prior depth maps from the acquired RGB images for supervised 3D scene reconstruction. In addition, we use the Gaussian encoding that approximates the conic truncation corresponding to the pixel as input to MLPs [35] to allow the system to adapt and capture spatial information more precisely. Moreover, we present reparameterization, which enables our SLAM system to achieve ideal reconstruction results in unbounded scenes. Finally, We introduce loop closure detection strategy for global pose adjustment to mitigate the scale drift. Extensive evaluations on a wide range of indoor RGB sequences demonstrate the scalability and predictive power of our approach. Overall, we make the following contributions:
@@
- We propose a novel NeRF-based dense SLAM system that can handle RGB input conditions and allow high-quality novel view synthesis by estimating the depth prior.
- We introduce a ray reparameterization strategy and a Gaussian encoding input that represents the region over which the radiance field should be integrated for neural fields to efficiently represent unbounded scenes.
- We demonstrate state-of-the-art performance on scene reconstruction based on multiple datasets and show competitive results with recent SLAM systems.
@@ 2 Related work
NeRF-based SLAM Recent advances in the Neural Radiance Field have shown powerful ability in scene reconstruction. In this evolutionary sequence, NICE-SLAM [46] and iMAP [30] were the first systems to incorporate these concepts, effectively employing NeRF [15] to render intricate environmental geometry by optimizing neural fields. While they established a strong foundation, the lack of robust loop closure detection became apparent. This results in a susceptibility to drift over time, especially in longer sequences or larger environments where re-localization is critical. GO-SLAM [43] introduces loop closure detection based on the previous work. It determines loop closure by detecting co-visibility between the current and historical keyframes and also attempts to map by relying only on RGB inputs. It has made significant progress in solving the scale drift problem, but the lack of the original depth information of the scene in its RGB input mode leads to a large error in scene scale reconstruction. Subsequently, HI-SLAM [42] emerged, striving to overcome the flaws inherent in GO-SLAM by implementing a more efficient loop closure and a priori depth estimates from Hessian matrices. However, the systems mentioned above are incapable of mapping in unbounded scenes.
Figure 2: Overview of MoD-SLAM. We demonstrate the monocular mode of our system, which uses RGB streams as its input. 1) Within the tracking process, the system ascertains the camera poses and depth values while executing loop closure based on a co-visibility check. Simultaneously, it selects key frames that are subsequently fed into the mapping module. 2) The mapping process leverages the RGB map derived from these keyframes to forecast depth values via a depth estimation module and depth distillation process. Concurrently, Gaussian encoding is applied to the keyframes, and the resultant spatial mean and covariance are fed into MLPs.
@@ 3 Method
Fig. 2 shows an overview of our method. For an input RGB stream, we employ the following two-step process: (1) we estimate the depth values of key frames using a depth estimator, and (2), we refine these depth estimates using depth distillation to filter inaccuracies (Sec. 3.1). In the tracking process, we use the obtained depth information $\left\{\mathbf{D}_{i}\right\}_{i=1}^{M}$ for pose optimization. And we also introduce the loop closure detection to correct the drift (Sec. 3.3). In the mapping process, we use Gaussian sampling to capture geometric structure and visual appearance information in the space. We also use the contraction function to deal with long-range spatial information in the borderless scene (Sec. 3.2). Finally, employing a NeRF-based volume rendering method (Sec. 3.4), we generate color and depth values for each pixel, facilitating system network optimization (Sec. 3.5).
@@ 3.1 Prior Depth Estimation
Our proposed monocular SLAM system contains a depth estimation module and a depth distillation module. More precisely, the input RGB stream serves dual purposes: first, as a dataset for normal tracking and mapping process, and second, as input to a pre-trained depth estimation module. This module generates depth information for the scene, which is then distilled through a depth distillation module to refine the depth data. The use of priori depth estimation significantly enhances the reconstruction accuracy.
@@
The findings from [3, 13] indicate that utilizing a singular metric head across various datasets fails to yield effective predictions. Moreover, we find that predicting depth using only the pre-trained ZoeDepth model leads to scale non-conformity. Therefore, we set multiple metric heads in our model and fine-tune ZoeDepth using specific datasets. These strategies constrain the scale of the depth information predicted by our depth estimation module.
@@
3.2 Scene Representation
Multivariate Gaussian Encoding  NeRF constructs the training dataset by projecting a ray to the center of each pixel point and then performing coarse and fine sampling. This approach leads to missing information. Inspired by the Mip-NeRF [1], we direct a cone-shaped projection to the pixel rather than slender rays. Projecting the cone allows us to collect more detailed information in the three-dimensional space and transfer it to MLPs for training. The training speed of the model is also greatly improved by eliminating the coarse and fine sampling processes.
To compute the above features, we approximate the conic truncation with a multivariate Gaussian to compute the features in a particular space. We introduce $t_{\mu}=(t_{0}+t_{1})/2$ and $t_{\delta}=(t_{0}-t_{1})/2$ as refined parameters to replace neighboring sampling points $t_{0}$ and $t_{1}$ to bolster the system’s stability. The multivariate Gaussian features are represented by three variables: the mean distance along the ray $\mu_{t}$, the variance along the ray $\sigma_{t}^{2}$, and the variance perpendicular to the ray $\sigma_{r}^{2}$:
following the above equations, we can obtain a multivariate Gaussian representation for the position as well as the direction in a particular space.
Figure 3: Scene Reparameterization. To manage unbounded scenes, we use a contract function to map mean and covariance data from spatial sampling (black dashed line) directly to a new range (red solid line). We save the data within the sphere of radius 1 (yellow region) while mapping the data from regions beyond a radius of 1 into the spherical coordinate system with a radius of 2 (orange region).
This can be interpreted as an extension of the normalized device coordinate (NDC). We make the scene shrink along the inverse depth direction, thus enabling the network to train on and interpret unbounded geometric scenes. However, since in Equ. 5 above we used the mean and covariance to represent information about the cross-section within the conic, we also need to map the mean and covariance to it.
where $\mathbf{J}_{f}(\mathbf{\mu})$ is the jacobian of $f$ at $\mathbf{\mu}$. After doing that, we can obtain mean $E(\mathbf{x})$ and covariance $D(\mathbf{x})$ with our choice for $f$ based on Equ. 6:
3.3 Loop Closure
3.4 Depth and Color Rendering
Inspired by the recent work in NeRF, we propose to also use Neural radiance field to reconstruct 3D scenes.
@@ 3.4 Depth and Color Rendering
Given the current camera pose and camera intrinsic matrix, we can calculate the viewing direction r and change it to the world coordinate. Unlike NeRF, we propose here a more efficient sampling technique, the differential sampling technique. To accomplish this idea, we first sample the ray at the proximal $n$ as well as at the distal $f$, and then obtain a series of sampling points by sampling uniformly in this space $\{s_{1},...,s_{N}\}$. After that, we map the sampling points to the following function:
@@
To complete the scene reconstruction, we randomly extract $N$ rays and $M$ pixel points in the current frame and key frames. We input the position set $\mathbf{P}$ to the depth prediction module in Sec. 3.1 to get the true depth $\mathbf{D}$. We do a multivariate Gaussian variation of the pose set $\mathbf{G}$ in Sec. 3.2 to get the mean as well as the covariance set. The mean and covariance are then passing into the MLPs to obtain the predicted volume density as well as the predicted color. Volume density is used to calculate the occupancy probability $o_{i}$ and the ray termination probability at each sample can then be calculated as $w_{i}=o_{i}\prod_{i=j}^{i-1}(1-o_{j})$. Eventually, the depth and color can be rendered as:
@@
3.5 Tracking and Mapping
Camera Tracking For more accurate position tracking, in this work, we employ DROID-SLAM [34] as the front end of our system due to its superior position prediction capabilities and minimal position drift. We take the RGB stream as input, and for each frame maintain two state variables: camera position and inverse depth. The pose set $\left\{\mathbf{G}_{t}\right\}^{N}_{t=0}$ and the inverse depth set $\left\{\mathbf{d}_{t}\right\}^{N}_{t=0}$ are unknown state variables, which are iteratively updated during the inference process of processing a new frame. Each input image is processed through a feature extraction network with a context network.
In the network training phase, we train our network by using a combination of pose loss and depth loss. We pre-compute a distance matrix of size $\mathbf{K}*\mathbf{K}$, where $\mathbf{K}$ is the length of each input image. This matrix is used to store the size of the average optical flow between each pair of frames. Then, we compute the depth loss $\mathcal{L}_{d}$ generated by the predicted depth and the true depth, and the position pose loss $\mathcal{L}_{v}$ generated by the predicted bit pose and the true position pose. The loss is taken as the average l2 distance between the two optical flow fields [41]:
$\centering\mathcal{L}_{tracking}=\mathcal{L}_{v}+\lambda_{v}\mathcal{L}_{d},\@add@centering$ (11)
where $\lambda_{v}$ is a weight used to limit the depth loss to the overall tracking loss. In this system, we take $\lambda_{v}$ equals to 0.1.
Mapping To enhance the fidelity of the reconstructed scene to the actual scene, we initially sample $M$ pixels from the keyframes. Subsequently, the network can be optimized by minimizing the loss.
@@ 3.5 Tracking and Mapping
We also use Equ. 1 and Equ. 2 to constrain depth value to supervise the reconstruction of this system and better recover the scale:
@@
$\centering\mathcal{L}_{dist}=\sum_{i,j}w_{i}w_{j}\lvert\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+1}}{2}\rvert+\sum_{i}\frac{1}{3}w_{i}^{2}(s_{i+1}-s{i}),\@add@centering$ (14)
@@
$\centering\mathcal{L}_{eik}=\frac{1}{MN_{ray}}\sum_{m=1}^{M}\sum_{i\in N}(1-\lVert\bigtriangledown\mathbf{\phi(s_{i})}\rVert)^{2}.\@add@centering$ (15)
@@
$\centering\mathcal{L}_{sdf}=\frac{1}{MN_{ray}}\sum_{m=1}^{M}\sum_{i\in N}\lvert\mathbf{\phi(s_{i})}-\mathbf{b(s_{i})}\rvert,\@add@centering$ (16)
@@ 4 Experiments
We evaluate our models in both synthetic and real-world datasets, including Replica [28], ScanNet [7] and TUM RGB-D [29]. We also conducted a comparative analysis of the reconstruction quality of our proposed system against state-of-the-art SLAM systems.

### 08 Mar 2024 v5

@@ Abstract
Monocular SLAM has received a lot of attention due to its simple RGB inputs and the lifting of complex sensor constraints. However, existing monocular SLAM systems are designed for bounded scenes, restricting the applicability of SLAM systems. To address this limitation, we propose MoD-SLAM, the first monocular NeRF-based dense mapping method that allows 3D reconstruction in real-time in unbounded scenes. Specifically, we introduce a Gaussian-based unbounded scene representation approach to solve the challenge of mapping scenes without boundaries. This strategy is essential to extend the SLAM application. Moreover, a depth estimation module in the front-end is designed to extract accurate priori depth values to supervise mapping and tracking processes. By introducing a robust depth loss term into the tracking process, our SLAM system achieves more precise pose estimation in large-scale scenes. Our experiments on two standard datasets show that MoD-SLAM achieves competitive performance, improving the accuracy of the 3D reconstruction and localization by up to 30% and 15% respectively compared with existing state-of-the-art monocular SLAM systems.
@@
3D Reconstruction Neural Radiance Field SLAM Monocular Depth Estimation Unbounded Scene
@@ 1 Introduction
In recent years, due to the development and advancement of Neural Radiance Field (NeRF) [15, 8] and view synthesis [19, 26, 1] in 3D scene reconstruction, dense visual SLAM mapping has become feasible. Early NeRF-based SLAM systems such as iMAP [30] and NICE-SLAM [45] have achieved ideal results in dense mapping and tracking. However, they all have the problems of decreasing accuracy and scale drift in large scenes. Subsequent researches [41, 42, 37, 6, 25] apply the signed distance function (SDF) [22] to reconstruct the 3D scene more accurately. Nevertheless, these approaches still present limited capability in unbounded scenes.
To represent unbounded scenes, a commonly adopted strategy is to employ spatial warping methods that map unbounded space to bounded space. However, warping implies information compression, which presents a significant challenge to the accuracy of features, especially for monocular SLAM systems. NEWTON [14] is the first RGB-D NeRF-based SLAM system capable of mapping unbounded scenes. Nevertheless, it lacks support for RGB input and relies on Nerfstudio [32] in back-end processing to achieve reconstruction of unbounded scenes. Thus, the applicability of this system is significantly limited.
@@
Figure 1: Reconstruction results on ScanNet 0207 [7]. We present MoD-SLAM, a neural-based monocular dense mapping method. Our method shows a more powerful capability than existing state-of-the-art work: GO-SLAM [42].
To address the above problems, we propose MoD-SLAM, the first monocular NeRF-based dense mapping method that allows 3D reconstruction in real-time in unbounded scenes. In MoD-SLAM, we introduce a novel reparameterization approach. This method contracts space in spherical coordinates, thus eliminating the NDC [15] forward-facing scene limitations. Furthermore, unlike the sampling approach of previous NeRF-based SLAM systems, we adopt Gaussian encoding to sample information by projecting a cone towards the pixel center. This method allows for a more accurate and comprehensive capture of spatial information. To improve the accuracy of reconstruction in unbounded scenes, we further incorporate a monocular depth estimation module and a depth distillation module at the front-end. Compared to previous monocular systems [25, 41] that directly invert depth from the Hessian matrix, our pre-trained depth model is allowed to fine-tuned on each scene. This crucial improvement enable MoD-SLAM achieves excellent predictive performance in various scenes, a feat unprecedented in prior work. Additionally, given the challenge of pose estimation in monocular SLAM systems, particularly in unbounded scenes, we introduce a depth-supervised camera tracking method. More specifically, we add a robust depth loss term to constrain scales, enabling the system to estimate poses more accurately. Extensive evaluations on a wide range of indoor RGB sequences demonstrate the scalability and predictive power of our approach. Overall, we make the following contributions:
@@
- We propose the first monocular NeRF-based dense SLAM system that allow unbounded scene representation and achieve high-quality novel view synthesis by estimating the depth values prior.
- We introduce a ray reparameterization strategy and a Gaussian encoding input that contract the scene and capture spatial features more accurately to efficiently represent unbounded scenes.
- We introduce a depth-supervised camera tracking method to improve camera pose estimation in monocular SLAM. By combining loop closure detection and global optimization, we demonstrate state-of-the-art performance in both mapping and tracking metrics of our system.
@@ 2 Related work
NeRF-based SLAM Recent advances in the Neural Radiance Field have shown powerful ability in scene reconstruction. In this evolutionary sequence, NICE-SLAM [45] and iMAP [30] were the first systems to incorporate these concepts, effectively employing NeRF [15] to render intricate environmental geometry by optimizing neural fields. However, the lack of loop closure detection results in a susceptibility to drift over time, especially in longer sequences or larger environments where re-localization is critical. GO-SLAM [42] introduces loop closure detection based on the previous work. It determines loop closure by detecting co-visibility between the current and historical keyframes and also attempts to map by relying only on RGB inputs. It has made significant progress in solving the scale drift problem, but the lack of the original depth information of the scene in its RGB input mode leads to a large error in scene scale reconstruction. Subsequently, HI-SLAM [41] emerged, striving to overcome the flaws inherent in GO-SLAM [42] by implementing a more efficient loop closure and a priori depth estimates from Hessian matrices. However, the systems mentioned above are incapable of mapping in unbounded scenes.
Figure 2: Overview of MoD-SLAM. We demonstrate the monocular mode of our system, which uses RGB streams as its input. 1) Within the tracking process, the system performs depth-supervised camera pose prediction and executes loop closure and global optimization based on a co-visibility check to refine camera poses and correct drift. Simultaneously, it selects keyframes to fed into the mapping module. 2) The mapping process leverages the RGB map derived from keyframes to forecast depth values via a depth estimation and depth distillation modules. Concurrently, Gaussian encoding and contraction function are applied to the keyframes, and the resultant spatial mean and covariance are fed into MLPs to reconstruct unbounded scenes.
@@ 3 Method
Fig. 2 shows an overview of our method. We aim to estimate the camera poses $\left\{\mathbf{R}_{i},\mathbf{T}_{i}\right\}_{i=1}^{M}$ of every frame and simultaneously reconstruct the unbounded scene with RGB streams input. In Sec. 3.1, we introduce the spherical contraction function to handle unbounded large scenes and use Gaussian encoding for a more precise geometric structure and visual appearance information $\left\{{\sigma}_{i},\mathbf{\Sigma}_{i}\right\}_{i=1}^{M}$ capture. To address the scale inconsistency of monocular unbounded scene reconstruction, in Sec. 3.2, we design a monocular depth estimation module and depth distillation module to extract accurate depth values $\left\{\mathbf{D}_{i}\right\}_{i=1}^{M}$, constraining the scale in the currently observed scene. In the camera tracking process, Sec. 3.3 introduces a robust depth penalty term to significantly improve pose prediction accuracy based on depth values generated by the depth estimation module. Sec 3.4 introduces a NeRF-based volume rendering method, which uses color values $\bigl{\{}\hat{\mathbf{C}}_{i}\bigr{\}}_{i=1}^{M}$ and depth values $\bigl{\{}\hat{\mathbf{D}}_{i}\bigr{\}}_{i=1}^{M}$ for networks training.
3.1 Gaussian Scene Representation
Multivariate Gaussian Encoding  NeRF constructs the training dataset by projecting a ray to the center of each pixel point and then performing coarse to fine sampling. This approach leads to missing information as rays only capture information along extension lines of pixel centers rather than on three-dimensional spaces. Inspired by the Mip-NeRF [1], we direct a cone-shaped projection to the pixel rather than slender rays. Projecting the cone allows us to collect more detailed information in the three-dimensional space and transfer it to MLPs for training. The training speed of the model is also greatly improved by eliminating the coarse to fine sampling process.
To compute the above features, we approximate the conic truncation with multivariate Gaussians to compute features in a particular space. We introduce $t_{\mu}=(t_{0}+t_{1})/2$ and $t_{\delta}=(t_{0}-t_{1})/2$ as refined parameters to replace neighboring sampling points $t_{0}$ and $t_{1}$ to bolster the system’s stability. The multivariate Gaussian features are represented by three variables: the mean distance along the ray $\mu_{t}$, the variance along the ray $\sigma_{t}^{2}$, and the variance perpendicular to the ray $\sigma_{r}^{2}$:
following the above equations, we can obtain a multivariate Gaussian representation for positions as well as directions in a particular space.
Figure 3: (a) Scene Reparameterization. To manage unbounded scenes, we use a contract function to map mean and covariance data from spatial sampling (black dashed line) directly to a new range (red solid line). We save the data within the sphere of radius 1 (yellow region) while mapping the data from regions beyond a radius of 1 into the spherical coordinate system with a radius of 2 (orange region). (b) NDC vs Our Reparameterization. NDC results in voids at the boundaries of geometric objects and lacks the feature extraction and reconstruction quality achieved by our reparameterization approach.
This can be interpreted as an extension of the normalized device coordinate (NDC). We make the scene shrink along the inverse depth direction, thus enabling the network to train on and interpret unbounded geometric scenes. Since in Equ. 3 above we used the mean and covariance to represent information about the cross-section within the conic, we also need to map the mean and covariance to it.
where $\mathbf{J}_{f}(\mathbf{\mu})$ is the jacobian of $f$ at $\mathbf{\mu}$. After doing that, we can obtain mean $E(\mathbf{x})$ and covariance $D(\mathbf{x})$ with our choice for $f$ based on Equ. 4:
Fig. 3(a) presents the visualization of our reparameterization. Compared to NDC reparameterization, our method solves the issue of missing scene information caused by the misalignment of the camera viewpoint with the primary object. Fig. 3(b) demonstrates that our method achieves better hole-filling result at the boundaries of geometric objects than NDC.
@@ 3.2 Prior Depth Estimation
The absence of a predefined scale in unbounded scenes poses a significant challenge to monocular reconstruction, leading to pronounced scale disparity. To improve the accuracy of our monocular implicit neural representation in unbounded scenes, we incorporate prior depth estimation for scale constraints. Our proposed monocular SLAM system contains a depth estimation module and a depth distillation module. More precisely, the input RGB stream serves dual purposes: first, as a dataset for normal tracking and mapping process, and second, as input to a pre-trained depth estimation module. This module generates depth information for the scene, which is then distilled through a depth distillation module to refine the depth value. With fine-tuning pre-trained depth estimation module, compared to previous monocular depth estimation methods, our method achieve more accurate depth estimation in various scenes.
@@
The findings from [3, 13] indicate that utilizing a singular metric head across various datasets fails to yield effective predictions. Moreover, we find that predicting depth using only the pre-trained ZoeDepth model leads to scale non-conformity. Therefore, we set multiple metric heads in our model and fine-tune ZoeDepth using specific datasets. These strategies constrain the scale of the depth values predicted by our depth estimation module.
@@
3.3 Depth-supervised Camera Tracking
For more accurate position tracking, in this work, we employ DROID-SLAM [34] as the front-end of our system. In traditional DROID-SLAM monocular mode of tracking, depth values are obtained by reverse-solving the Hessian matrix from the predicted poses of the GRU layer. The cost function as follow:
$\centering\mathbf{E}(\mathbf{G^{\prime}},\mathbf{D^{\prime}})=\sum_{(i,j)\in\epsilon}\|\mathbf{p}_{ij}^{*}-\delta(\mathbf{p}_{i},\mathbf{d}^{\prime}_{i})\|_{diag(w_{ij})}^{2},\@add@centering$ (9)
where $\|\cdot\|_{diag()}$ is the Mahalanobis distance which weights the error terms based on the confidence weights $w_{ij}$. $\mathbf{p}_{ij}^{*}$ represents the corrected coordinates of pixel $\mathbf{p}_{i}$ mapped into frame $j$ and $\delta(\cdot)$ denotes the coordinates of pixel $\mathbf{p}_{i}$ mappend into frame $j$ using the refined depth. However, the depth values are subject to scale uncertainty because of the error in the pose prediction. To address this issue, we constrain depth values by introducing a new depth penalty term into the original loss function as follow:
$\centering R=\sum_{i\in\epsilon}\begin{cases}(\mathbf{d}_{i}-\mathbf{d}^{\prime}_{i})^{2}&for|\mathbf{d}_{i}-\mathbf{d}^{\prime}_{i}|<\tau_{tra}\\ \tau_{tra}|\mathbf{d}_{i}-\mathbf{d}^{\prime}_{i}|&otherwise\end{cases},\@add@centering$ (10)
where $\mathbf{d}_{i}$ denotes the depth value obtained from the depth estimation module (Sec. 3.2) and $\mathbf{d}^{\prime}_{i}$ represents the depth value from Hessian matrix. $\tau_{tra}$ is the hyperparameter. Due to the error of the depth values produced by the depth estimation module, we utilize the L1 loss for outlier pixels, while employing the L2 loss for the rest. Compared to solely employing the L2 loss, our penalty term is more robust to outliers. Then, following classical optical flow-based SLAM methods [40], we use the Gauss-Newton method to solve the Hessian matrix for obtaining updates of inverse depth and camera poses:
$\left[\begin{array}[]{cc}C&E\\ E^{T}&P\\ \end{array}\right]\left[\begin{array}[]{c}\mathbf{\Delta}\mathbf{\xi}\\ \mathbf{\Delta}\mathbf{d}\\ \end{array}\right]=\left[\begin{array}[]{c}\mathbf{v}\\ \mathbf{w}\\ \end{array}\right]\qquad\begin{aligned} \mathbf{\Delta\xi}&=[C-EP^{-1}E^{T}]^{-1}(\mathbf{v}-EP^{-1}\mathbf{w})\\ \mathbf{\Delta d}&=P^{-1}(w-E^{T}\mathbf{\Delta\xi})\end{aligned}$ (11)
where $C$ is the block camera matrix, and $P$ is the diagonal matrix corresponding to the points. $\mathbf{\Delta\xi}$ and $\mathbf{\Delta d}$ denotes pose and depth update values. Benefiting from prior depth values, our system can provide a more precise estimate of the pose. We also provide details on loop closure detection in Appendix.
3.4 Mapping
@@
Depth and Color Rendering Given the current camera pose and camera intrinsic matrix, we can calculate the viewing direction r and change it to the world coordinate. Unlike NeRF, we propose here a more efficient sampling technique, the differential sampling technique. To accomplish this idea, we first sample the ray at the proximal $n$ as well as at the distal $f$, and then obtain a series of sampling points by sampling uniformly in this space $\{s_{1},...,s_{N}\}$. After that, we map the sampling points to the following function:
@@
To complete the scene reconstruction, we randomly extract $N$ rays and $M$ pixel points in the current frame and key frames. We input the position set $\mathbf{P}$ to the depth prediction module in Sec. 3.2 to get the true depth $\mathbf{D}$. We do a multivariate Gaussian variation of the pose set $\left\{\mathbf{R},\mathbf{T}\right\}$ in Sec. 3.1 to get the mean as well as the covariance set. The mean and covariance are then passing into the MLPs to obtain the predicted volume density as well as the predicted color. Volume density is used to calculate the occupancy probability $o_{i}$ and the ray termination probability at each sample can then be calculated as $w_{i}=o_{i}\prod_{i=j}^{i-1}(1-o_{j})$. Eventually, the depth and color can be rendered as:
@@
Loss To enhance the fidelity of the reconstructed scene to the actual scene, we initially sample $M$ pixels from the keyframes. Subsequently, the network can be optimized by minimizing the loss.
@@
We also use Equ. 7 and Equ. 8 to constrain depth value to supervise the reconstruction of this system and better recover the scale:
@@
$\centering\mathcal{L}_{dist}=\sum_{i,j}w_{i}w_{j}\lvert\frac{s_{i}+s_{i+1}}{2}-\frac{s_{j}+s_{j+1}}{2}\rvert+\sum_{i}\frac{1}{3}w_{i}^{2}(s_{i+1}-s_{i}),\@add@centering$ (16)
@@
$\centering\mathcal{L}_{eik}=\frac{1}{MN_{ray}}\sum_{m=1}^{M}\sum_{i\in N}(1-\lVert\bigtriangledown\phi(\mathbf{s}_{i})\rVert)^{2}.\@add@centering$ (17)
@@
$\centering\mathcal{L}_{sdf}=\frac{1}{MN_{ray}}\sum_{m=1}^{M}\sum_{i\in N}\lvert\phi(\mathbf{s}_{i})-\mathbf{b}(\mathbf{s}_{i})\rvert,\@add@centering$ (18)
@@
4 Experiments
We evaluate our models in both synthetic and real-world datasets, including Replica [28] and ScanNet [7]. We conducted a comparative analysis of the reconstruction quality of our proposed system against state-of-the-art SLAM systems. We also show more experimental results in the Appendix.